OpenStudy (luigi0210):

Integration:

OpenStudy (luigi0210):

\[\LARGE \int_{0}^{\frac{T}{2}}(sin\frac{2\pi tT}{T}-\alpha)~dt \]

OpenStudy (dan815):

im here luigi my boy

OpenStudy (luigi0210):

I think we use u-sub, right?

OpenStudy (dan815):

if T is just another constat

OpenStudy (dan815):

yeah you can do u sub, but for these kind learn to do it just straight

OpenStudy (dan815):

you can see that derivative of -cos = sin

OpenStudy (luigi0210):

I got that far.. it's the alpha that's throwing me off.

OpenStudy (anonymous):

\[\int\limits \sin(2pit)dt-\alpha \int\limits 1 dt\]

OpenStudy (dan815):